OpenStudy (anonymous):

Use a finite sum to estimate the average value of the function on the given interval by partitioning the interval and evaluating the function at the midpoints of the subintervals. f(x) = 3x^5 on [1, 3] divided into 4 subintervals

OpenStudy (anonymous):

your interval has length 2, so if you divide it in to four parts they will have length \(\frac{1}{2}\)

OpenStudy (anonymous):

the intervals will have endpoints at \[1,1.5,2,2.5,3\] and if you go half way between each the midpoints will be at \[1.25,1.75,1.25,2.75\]

OpenStudy (anonymous):

your job is therefore to evaluate the function at each of these points, add up the total, and then divide the result by the length of the interval, which is 2

OpenStudy (anonymous):

how do i evaluate it

OpenStudy (anonymous):

\[f(x) = 3x^5 \] \[f(1.25)=4(1.25)^5\] etc use a calculator

OpenStudy (anonymous):

added them up and divided by 2. i assumed that 4(1.25)^5 was a typo. i used 3(x)^5. so after dividing by 2 i got 270.299804688